OpenStudy (anonymous):

Rewrite f(x) = 2(x − 1)2 + 3 from vertex form to standard form.

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

actually this is the new question

OpenStudy (anonymous):

Rewrite f(x) = 3(x − 2)2 + 1 from vertex form to standard form.

OpenStudy (anonymous):

@fred7

OpenStudy (anonymous):

thank you so much the smartone

OpenStudy (anonymous):

(1,3) is your vertex

OpenStudy (anonymous):

wait so how does that help me

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

This is in General form\[2(x-1)^2+3\]

OpenStudy (anonymous):

aight

OpenStudy (anonymous):

It is easy to find the vertex in general form , by equaling (x-1)=0

OpenStudy (anonymous):

wait which one is this the first or second one cause i dont need the first one any more

OpenStudy (anonymous):

From that you get (1,y)

OpenStudy (anonymous):

Now you need to find the y value

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

The y value is always the constant

OpenStudy (anonymous):

so tell whats the constant

OpenStudy (anonymous):

@braaaahhh

OpenStudy (anonymous):

srry um im still lost is this the first or second question

OpenStudy (anonymous):

wats a constant @fred7

OpenStudy (anonymous):

A constant is a number that does not very due to a variable. The constant in 3x+9=y is 9,because it does not have a variable

OpenStudy (anonymous):

so the constant for two would be three

OpenStudy (anonymous):

@fred7

OpenStudy (anonymous):

YA

OpenStudy (anonymous):

sO THEN (1,y) when why is the constant so whats the vertex

OpenStudy (anonymous):

So then the vertex is (1,3)

OpenStudy (anonymous):

You got X from (x-1)=0 and you got Y by finding the constant. Do you get it now

OpenStudy (anonymous):

Yes tahnk i got the answer on the test right thank you